OpenStudy (anonymous):

Find the equation of the line that passes through the point (-3, -8) with a slope of -5. Write your answer in Slope-Intercept Form.

Parth (parthkohli):

\[y - y_1 = m(x - x_1)\]Here, \(m = -5\), \(x_1 = -3\) \(x_2 = -8\)

Parth (parthkohli):

After you get your answer, rearrange it in the form of \(y = mx + b\)

Parth (parthkohli):

\[x_1 = -3\]\[y_1 = -8\]*

OpenStudy (anonymous):

y−y1=m(x−x1) y + 8 = -5 ( x + 3) y + 8 = -5x - 15 y = -5x - 23